OpenStudy (anonymous):

Martin’s Hardware is clearing out their inventory of bicycles and tricycles. The clerk rolls all of the bicycles and tricycles outside and arranges them in a line on the sidewalk. If the total number of bicycles and tricycles is 37 and the total number of wheels on all the bicycles and tricycles combined is 90, how many tricycles are in the line on the sidewalk? For this question, wouldn't you use linear programming? If so, can someone explain and help me with it? I really don't understand.

OpenStudy (phi):

I think this is just ordinary algebra. The first step is assign variable names to the "unknowns"

OpenStudy (anonymous):

Let b and t be the number of bicycles and tricycles respectivley. Solve the following equations for b and t:\[\{b+t=37,2b+3t=90\}\]\[\{b=21,t=16\} \]

OpenStudy (phi):

they ask for how many tricycles so T for # of tricycles would be a good name for a variable they say total number of bicycles and tricycles is 37 they mention the number of bicycles and tricycles. so B for # of bicycles would be good. translate the sentence total number of bicycles and tricycles is 37 into math B + T = 37 that is short for "# of bikes plus # of trikes is 37

OpenStudy (phi):

you also have to translate total number of wheels on all the bicycles and tricycles combined is 90 into math. tobey already posted the equations, but if you want to figure it out yourself, you have to think" how many wheels does a bike have?" and then " if I have B bikes (B is the (unknown) # of bikes) ) how many wheels is that? " you should think: 2 wheels per bike times B bikes: 2B wheels use the same idea for the number of wheels on a trike to get 3T wheels add them up to get 2B+3T = 90

OpenStudy (phi):

once you have the equations (remember: for every unknown you need an equation... in this case you have two unknowns, B and T, and 2 equations) so you can solve for both variables. B + T = 37 2B+3T = 90 can you solve these?

OpenStudy (anonymous):

Thank you!!

OpenStudy (phi):

just to finish, you can solve using elimination or substitution elimination: multiply the first equation by -2 (that means both sides, all terms) to get -2B + -2T = -2*37 add this equation to the second equation 2B+3T = 90 to get -2B + 2B + -2T + 3T = -2*37 + 90 and simplify

OpenStudy (phi):

-2B + 2B is 0 and -2T + 3T is 1T, so the left side is 1T = -2*37 + 90 or T = -74 + 90 or T= 90 - 74 = 16